Readings: Review for binomial distributions. 6. A plan for an executive travelers club has been developed by an airline on the premise that 10% of its current customers would qualify for membership. Assume the premise is valid. (a) What is the chance that between 2 and 4 (inclusive) customers out of 25 randomly selected current customers qualify for membership? (b) What is the expected number of customers who qualify and the standard deviation of the number who qualify in a random sample of 100 current customers?
Answer #1

Ans:

a)n=25 and p=0.1

P(2<=x<=4)=P(x=2)+P(x=3)+P(x=4)

=25C2*0.12*0.923+25C3*0.13*0.922+25C4*0.14*0.921

=0.2659+0.2265+0.1384

=0.6308

b)

n=100 and p=0.1

mean=np=100*0.1=10

standard deviation=sqrt(np(1-p))=sqrt(100*0.1*0.9)=3